# A Comprehensive Survey of Direct Preference Optimization: Datasets, Theories, Variants, and Applications

### v2

@@ 4.2 AI labeled
Davanstrien/dataset-tldr-preference-dpo. The davanstrien/dataset-tldr-preference-dpo dataset (davanstrien, 2024) is designed for summarization and text generation tasks using DPO. It was constructed by creating concise summaries (tl;dr) (Völske et al., 2017) of dataset cards, specifically aiming to highlight the most critical aspects of each dataset. These summaries were generated using LLMs like Meta-Llama-3-70B-Instruct (AI@Meta, 2024) and Nous-Hermes-2-Mixtral-8x7B-DPO ("karan4d" "huemin_art", 2024). The dataset includes prompts instructing the model to generate brief, informative summaries based on the dataset cards. The goal is to make these summaries useful for quickly understanding the dataset without unnecessary information, such as dataset size or license details. The dataset was filtered to ensure that only examples with meaningful differences between chosen and rejected responses were included.
@@
Zerolink/Zsql-postgres-dpo.The zerolink/zsql-postgres-dpo dataset (Zerolink, 2024a) is designed for training models to convert natural language questions into SQL queries, specifically in the PostgreSQL dialect. This dataset contains around 200,000 pairs of DPO examples, making it ideal for text-to-SQL generation tasks.The construction of the dataset involves selecting SQL queries that are optimized for their syntactic and computational efficiency. For each example, the dataset includes a database schema, a natural language question, and two SQL queries: a "chosen" query (the preferred one) and a "rejected" query. These preferences are determined based on minimizing complexity and optimizing execution.
@@
RewardBench (Lambert et al., 2024) is a benchmark dataset evaluating reward model’s performance on various queries. Rewardbench is a collection of chosen or reject responses as well as prompts over tasks such as chatting, reasoning, and safety. In chat task, Rewardbench collects subsets from AplacalEval (Li et al., 2023b) including alpacaeval-easy, alpacaeval-length, alpacaeval-hard and MT Bench (Zheng et al., 2023a) subset including, mt-bench-easy, mt-bench-medium. In safety category, Rewardbench collects subsets like efusals-dangerous, refusals-offensive, xstest-should-refuse (Röttger et al., 2024), xstest-should-respond, do not answer (Wang et al., 2023d). In reasoning task, Rewardbench encompasses datasets like PRM Math (Lightman et al., 2023), HumanEvalPack (Muennighoff et al., 2024) on multiple programming languages. RewardBench also collects datasets from prior datasets including Anthropic Helpful, Stanford Human Preferences (SHP), and OpenAI’s Learning to Summarize. The construction of dataset is done by scoring chosen-rejected-prompt with scalar scores. The success situation is when the chosen response attains higher scores than the reject responses. Rewardbench contains 8108 rows of samples in all.
@@
MMInstruction/VLFeedback. VLFeedback (Li et al., 2023a) is a AI labeled dataset for large vision language models(LVLMs), aiming to enhance LVLMs’ ability to generate factual and helpful responses following the visual context. Given multi-modal instructions from different datasets, 12 LVLMs are prompted to generate responses, promising that each data sample in VLFeedback contains 4 responses from different models. Then, GPT-4V is employed to evaluate the quality of the generated responses based on the criteria of helpfulness, visual factuality, and ethical considerations. At last, the LLM-annotated preference ranking dataset consists more than 380k comparison pairs and 80k multimodal instructions.
@@
sqrti/SPA-VL. SPA-VL (Zhang et al., 2024a) is a safety-specific dataset for safety alignment of Vision language Models(VLMs). SPA-VL is a GPT-4V labeled alignment dataset consisting of 100,788 samples in a wide range including 6 harmfulness domains, 13 categories, and 53 subcategories. Each preference pair contains the given question and image as well as the chosen response and rejected response. The primary goal of this dataset is to improve the harmlessness and helpfulness of VLMs while preserving their core capabilities. In practice, all the images in this dataset are collected from the LAION-5B dataset (Schuhmann et al., 2022). The accompany questions are generated by Gemini 1.0 Pro Vision (Team et al., 2024). To ensure the diversity of answers, SPA-VL employs 12 different models to sample different responses. Preference annotations are conducted by GPT-4V in order to determine which response is preferred and which one is disprederred based on harmlessness and helpfulness.
@@
Stack-exchange-preferences (Lambert et al., 2023) is an open dataset containing questions and answers from Stack Overflow Data Dump. The purpose for this dataset is to enhance model’s preference training and instruction fine-tuning. The questions from the stack overflow have been filtered to fit the criteria aligned with preference models.
@@
Nectar (Zhu et al., 2023) is the first 7-wise preference dataset consisting of 182,954 rows of samples generated through GPT-4. Nectar dataset is composed of diverse chat prompts, high-quality and diverse responses, and accurate ranking labels. Prompts are sourced from many public dataset including lmsys-chat-1M (Zheng et al., 2023b), ShareGPT, Antropic/hh-rlhf (Bai et al., 2022a), UltraFeedback (Cui et al., 2023a), Evol-Instruct (Longpre et al., 2023), and Flan (Xu et al., 2023). Seven responses per prompt are primarily generated by a variety of models, including GPT-4, GPT-3.5-turbo, GPT-3.5-turbo-instruct.
@@
Orca (Lian et al., 2023) is subset of OpenOrca (Mukherjee et al., 2023) dataset. OpenOrca collects augmented FLAN data collection (Longpre et al., 2023), aiming at helping model to learning reasoning capability of the large founddation models through imitation learning. SlimOrca is a carefully curated subset of OpenOrca data, which apllys an extra pass by using GPT-4 to remove answers that are considered as wrong based on human annotations from FLAN dataset. This step raises the overall quality of Orca dataset and reduces dataset size, minimizing the computational requirements.
@@
Alpaca Farm GPT-4 (Dubois et al., 2024) is a result dataset from the simulation framework for learning from human feedback.  (Dubois et al., 2024) come up with a new simulator that enables model to generate human preference for scientific research at low cost. Based on the previous instruction dataset, Alpaca farm is firstly designing GPT-4 prompt to generate preference similar to human annotators. Variability is also considered in the procedure of the pipeline for simulating human annotation. After the auto-evaluation, Alpaca farm can produce high-quality data that contain preference feedback similar to human annotation.
@@ 5.1 Application on Large Language Models
Qwen2. At the fine-tuning stage, models within the Qwen2 (Yang et al., 2024d) series undergo supervised fine-tuning(SFT) as well as Direct Preference Optimization (DPO). Through this preference learning, the models are endowed with enhanced abilities to follow instructions, improving their performance across diverse tasks and making them more responsive to human users. The preference learning of qwen 2 models consists of two stages: offline training and online training Yang et al. (2024d). In the online training stage, the model further refines its performance by interacting with real-time feedback. During this phase, Qwen2 samples multiple candidate responses from the current policy, and a reward model evaluates and ranks them, selecting the most and least preferred responses to form preference pairs. These preference pairs are then fed back into the DPO process, allowing the model to iteratively improve its decision-making and instruction-following capabilities with each episode. This real-time refinement further aligns the model with human preferences, ensuring that it delivers more accurate and contextually appropriate responses. By integrating DPO at both stages, Qwen2 not only enhances its alignment with human preferences but also optimizes resource efficiency, making it suitable for handling large-scale models.
@@
The Llama 3 Herd of Models. In Llama3 Herd of Models (AI@Meta, 2024), post-training procedure includes relatively simple methods such as supervised fine-tuning(SFT), reject sampling(RS) and direct preference optimization(DPO) rather than complex reinforcement learning algorithms that induces extra complexity. After the supervised fine-tuning stage, several batches of preference data have been collected using the best performing models from the previous rounds of training. The distribution of the policy model is optimized in each traning round based on the collected quality data. The Llama3 team also confirms that compared to on-policy alogorithms like PPO (Schulman et al., 2017b), DPO does reduce computational costs and perform better, especially on instruction following benchmarks like IFEval (Zhou et al., 2023). To levarage the advantages of DPO, Llama3 team applies two modifications to the original DPO algorithm. Firstly, special formatting tokens including header and terminator are masked out in from both chosen and rejected responses in the loss to stabilize the DPO training, which is based on the observation that these special tokens contributing to loss may induce undesired model behaviors such as tail repetition or abruptly generating termination tokens. Secondly, an additional negative log-likelihood (NLL) loss term is applied to DPO with a scaling coefficient of 0.2 on the chosen sequences, helping further stabilize DPO by maintaining desired formatting for generation and preventing the decrease of log probability of chosen responses (Pang et al., 2024).Enhanced Safety DPO approach is a crucial part of ensuring that the model not only follows instructions accurately but also adheres to strict safety guidelines, reducing the likelihood of generating harmful or inappropriate content.
@@
DeepSeek LLM. DeekSeek LLM (DeepSeek-AI et al., 2024) base models are enhanced through post-training using DPO.  DeepSeek-AI et al. (2024) observed an increase in repetitive output as the proportion of math SFT data increased, which they attributed to the repetitive reasoning patterns present in such data. Weaker models often struggle to generalize these patterns effectively, leading to undesirable repetitive responses. To address this issue and enhance the models’ robustness, DPO was employed to further align the SFT models by utilizing a preference-based training approach. This approach leveraged pair-wise comparisons based on criteria such as helpfulness and harmlessness. Upon evaluation, the DPO-trained models demonstrated significant improvements across almost all key metrics, including performance in multilingual tasks such as creative writing, question answering, and instruction following. Moreover, the models also exhibited enhanced mathematical reasoning and logical deduction skills, underscoring the positive impact of the DPO training process in refining language model alignment.
@@
Mixtral of Experts. Mixtral 8x7B (Jiang et al., 2024a) is a Sparse Mixture of Experts (SMoE) model with publicly available weights. To improve its ability to follow complex instructions,  Jiang et al. (2024a) applied a two-step training approach. First, they employed Supervised Fine-Tuning (SFT) on an instruction-focused dataset to enhance model’s responsiveness to diverse commands. This was followed by DPO on a dataset with paired human feedback, allowing for further refinement based on user preferences. After completing the post-training process, Mixtral’s performance significantly exceeded that of well-known models such as GPT-3.5 Turbo (OpenAI, 2022), Claude-2.1 (Anthropic, 2024), Gemini Pro (Team et al., 2023), and Llama 2 70B (Touvron et al., 2023c) in human evaluation benchmarks, highlighting the effectiveness of this combined training methodology.
@@
Baichuan2. Baichuan 2 is a series of large-scale multilingual language models with 7 billion and 13 billion parameters, trained from scratch on a vast dataset of 2.6 trillion tokens (Yang et al., 2023). In the post-training stage, Baichuan2 is aligned with limited corpus of human-annotated data to improve the model’s performance with respect to vulnerability safety issues. To be specific,  (Yang et al., 2023) is algined with a red-teaming process that incorporates six types of safety attacks and over 100 granular safety value categories. Around 1K annotated data was used for initialization. By leveraging DPO, Baichuan 2 effectively addresses specific vulnerability issues, such as generating harmful or biased content, even with a limited amount of annotated safety-related data. By integrating DPO, Baichuan 2 demonstrates significant improvements in handling complex safety challenges, mitigating risks like the generation of misinformation, biased content, or other undesirable behaviors that can arise in large-scale multilingual models.
@@
Yi: Open Foundation Models. Yi model family (AI et al., 2024) encompasses a series of language and multi-modal models. The training of Yi model family proves that delicate fine-tuning framework can effectively save computational costs while maintaining same training results. In typical fine-tuning framework, reference or reward model is used to predict a batch of data or give a scalar reward value after certain action in specific environment. The time and computational resource spent on reference/reward model can not be neglected. In the next step, the target model will use data from previous step to calculate loss and carry out gradient descent in order to update parameters. In the training framework of Yi model family, multi-model scheduling framework is adopted to facilitate support multiple backends for different LLMs in a single job (AI et al., 2024). Compared to PPO, DPO algorithm can facilitate the intermediate results from the reference mode. Reusing data from previous stage effectively improves the training speed and save the resources costs.

### v3

@@ 4.2 AI labeled
Davanstrien/dataset-tldr-preference-dpo. The davanstrien/dataset-tldr-preference-dpo dataset davanstrien (2024) is designed for summarization and text generation tasks using DPO. It was constructed by creating concise summaries (tl;dr) Völske et al. (2017) of dataset cards, specifically aiming to highlight the most critical aspects of each dataset. These summaries were generated using LLMs like Meta-Llama-3-70B-Instruct AI@Meta (2024) and Nous-Hermes-2-Mixtral-8x7B-DPO "karan4d" "huemin_art" (2024). The dataset includes prompts instructing the model to generate brief, informative summaries based on the dataset cards. The goal is to make these summaries useful for quickly understanding the dataset without unnecessary information, such as dataset size or license details. The dataset was filtered to ensure that only examples with meaningful differences between chosen and rejected responses were included.
@@
Zerolink/Zsql-postgres-dpo.The zerolink/zsql-postgres-dpo dataset Zerolink (2024a) is designed for training models to convert natural language questions into SQL queries, specifically in the PostgreSQL dialect. This dataset contains around 200,000 pairs of DPO examples, making it ideal for text-to-SQL generation tasks.The construction of the dataset involves selecting SQL queries that are optimized for their syntactic and computational efficiency. For each example, the dataset includes a database schema, a natural language question, and two SQL queries: a "chosen" query (the preferred one) and a "rejected" query. These preferences are determined based on minimizing complexity and optimizing execution.
@@
RewardBench Lambert et al. (2024) is a benchmark dataset evaluating reward model’s performance on various queries. Rewardbench is a collection of chosen or reject responses as well as prompts over tasks such as chatting, reasoning, and safety. In chat task, Rewardbench collects subsets from AplacalEval Li et al. (2023b) including alpacaeval-easy, alpacaeval-length, alpacaeval-hard and MT Bench Zheng et al. (2023a) subset including, mt-bench-easy, mt-bench-medium. In safety category, Rewardbench collects subsets like efusals-dangerous, refusals-offensive, xstest-should-refuse Röttger et al. (2024), xstest-should-respond, do not answer Wang et al. (2023d). In reasoning task, Rewardbench encompasses datasets like PRM Math Lightman et al. (2023), HumanEvalPack Muennighoff et al. (2024) on multiple programming languages. RewardBench also collects datasets from prior datasets including Anthropic Helpful, Stanford Human Preferences (SHP), and OpenAI’s Learning to Summarize. The construction of dataset is done by scoring chosen-rejected-prompt with scalar scores. The success situation is when the chosen response attains higher scores than the reject responses. Rewardbench contains 8108 rows of samples in all.
RewardBench2 Malik et al. (2025), a new benchmark based on classification tasks, aims to provide challenging new data for RM evaluation based on accuracy. RewardBench2 uses new human prompts that have never been used for downstream evaluation (mostly from WildChat) as the main data source, promoting stricter evaluation and avoiding potential contamination with downstream evaluation goals. Its evaluation setting shifts from the common two-option (chosen and rejected) assessment to a four-option (one Chosen and three Rejected) assessment, reducing the random baseline from 50% to 25%, thus leaving more room for model improvement and enhancing score robustness, especially on more challenging subsets. RewardBench2 covers six domains, which are constructed through manual, programmatic, and LLM-based filtering techniques, aiming to improve the evaluation of challenging areas in existing RM benchmarks (Focus, Math, Safety) and add new challenging areas (Factuality, Precise Instruction Following, Ties).
VL-RewardBench Li et al. (2025b) is a comprehensive benchmark designed to evaluate the application of visual-language generation reward models (VL-GenRM) in visual perception, hallucination detection, and reasoning tasks. This benchmark consists of 1,247 high-quality examples, each carefully selected and composed of elements such as query, image, response, and human ranking. The queries are sourced from renowned datasets such as WildVision, RLAIF-V, POVID, and RLHF-V to ensure their quality. After screening by small models and human evaluation, challenging tasks were retained. Powerful commercial models generated responses and preference labels, and the final preference pairs were selected after human verification.
@@
MMInstruction/VLFeedback. VLFeedback Li et al. (2023a) is a AI labeled dataset for large vision language models(LVLMs), aiming to enhance LVLMs’ ability to generate factual and helpful responses following the visual context. Given multi-modal instructions from different datasets, 12 LVLMs are prompted to generate responses, promising that each data sample in VLFeedback contains 4 responses from different models. Then, GPT-4V is employed to evaluate the quality of the generated responses based on the criteria of helpfulness, visual factuality, and ethical considerations. At last, the LLM-annotated preference ranking dataset consists more than 380k comparison pairs and 80k multimodal instructions.
@@
sqrti/SPA-VL. SPA-VL Zhang et al. (2024a) is a safety-specific dataset for safety alignment of Vision language Models(VLMs). SPA-VL is a GPT-4V labeled alignment dataset consisting of 100,788 samples in a wide range including 6 harmfulness domains, 13 categories, and 53 subcategories. Each preference pair contains the given question and image as well as the chosen response and rejected response. The primary goal of this dataset is to improve the harmlessness and helpfulness of VLMs while preserving their core capabilities. In practice, all the images in this dataset are collected from the LAION-5B dataset Schuhmann et al. (2022). The accompany questions are generated by Gemini 1.0 Pro Vision Team et al. (2024). To ensure the diversity of answers, SPA-VL employs 12 different models to sample different responses. Preference annotations are conducted by GPT-4V in order to determine which response is preferred and which one is disprederred based on harmlessness and helpfulness.
@@
Stack-exchange-preferences Lambert et al. (2023) is an open dataset containing questions and answers from Stack Overflow Data Dump. The purpose for this dataset is to enhance model’s preference training and instruction fine-tuning. The questions from the stack overflow have been filtered to fit the criteria aligned with preference models.
@@
Nectar Zhu et al. (2023) is the first 7-wise preference dataset consisting of 182,954 rows of samples generated through GPT-4. Nectar dataset is composed of diverse chat prompts, high-quality and diverse responses, and accurate ranking labels. Prompts are sourced from many public dataset including lmsys-chat-1M Zheng et al. (2023b), ShareGPT, Antropic/hh-rlhf Bai et al. (2022a), UltraFeedback Cui et al. (2023a), Evol-Instruct Longpre et al. (2023), and Flan Xu et al. (2023). Seven responses per prompt are primarily generated by a variety of models, including GPT-4, GPT-3.5-turbo, GPT-3.5-turbo-instruct.
@@
Orca Lian et al. (2023) is subset of OpenOrca Mukherjee et al. (2023) dataset. OpenOrca collects augmented FLAN data collection Longpre et al. (2023), aiming at helping model to learning reasoning capability of the large founddation models through imitation learning. SlimOrca is a carefully curated subset of OpenOrca data, which apllys an extra pass by using GPT-4 to remove answers that are considered as wrong based on human annotations from FLAN dataset. This step raises the overall quality of Orca dataset and reduces dataset size, minimizing the computational requirements.
@@
Alpaca Farm GPT-4 Dubois et al. (2024) is a result dataset from the simulation framework for learning from human feedback.  Dubois et al. (2024) come up with a new simulator that enables model to generate human preference for scientific research at low cost. Based on the previous instruction dataset, Alpaca farm is firstly designing GPT-4 prompt to generate preference similar to human annotators. Variability is also considered in the procedure of the pipeline for simulating human annotation. After the auto-evaluation, Alpaca farm can produce high-quality data that contain preference feedback similar to human annotation.
CodeSteer-DPO-Dataset Chen et al. (2025) This dataset is derived from the training requirements of the CodeSteer model. This model is responsible for guiding large language models to solve problems correctly and efficiently by using the text/code approach, with particular emphasis on its ability in symbolic calculation. This dataset is played by GPT-4o as both the instructor and the responder. The basic dataset is generated through multiple rounds of dialogue and the final preference dataset is obtained through preference ranking. This dataset contains 4,462 guidance comparison pairs. Each guidance contains guidance of different degrees on the problem. These guidance include guidance on aspects such as the selection of text/code, algorithm usage, and algorithm emphasis.
ANAH Gu et al. (2025) ANAH is a bilingual dataset that offers analytical annotation of hallucinations in LLMs within generative question answering. It can be used to train hallucination detector and perform factulaity alignment.The dataset contains the literature corresponding to the topic. According to the questions selected from the literature, the answers from GPT3.5 and InternLM, as well as the analysis annotations for the responses of GPT3.5 and InternLM respectively.Each answer sentence in this dataset undergoes rigorous annotation, involving the retrieval of a reference fragment, the judgment of the hallucination type, and the correction of hallucinated content. ANAH consists of 12k sentence-level annotations for 4.3k LLM responses covering over 700 topics, constructed by a human-in-the-loop pipeline.
Unified Preference Dataset  Wang et al. (2025b) A comprehensive human preference dataset that spans multiple vision-related tasks. It integrate existing datasets and preprocess them to construct the first large-scale unified human preference dataset, which consists of approximately 236K data covering both image and video understanding and generation tasks.This dataset is a preprocessed and integrated version of eleven datasets, each of which contains preference rankings or ratings for images or videos.
MM-IFDPO-23k Construction Ding et al. (2025) This is a dataset used to enhance the Instruction Following ability of large models. The size of this dataset is 23k. Each row contains a conversation, an image, a selected answer and a rejected answer. It is used to help MLLMS better understand instructions and the responses users want. It defines 32 constraints to generate high-quality instruction-answer pairs. DPO training is conducted by generating rejection pairs by reducing the number of constraints.
@@ 5.1 Application on Large Language Models
Qwen2. At the fine-tuning stage, models within the Qwen2 Yang et al. (2024d) series undergo SFT as well as DPO. Through this preference learning, the models are endowed with enhanced abilities to follow instructions, improving their performance across diverse tasks and making them more responsive to human users. The preference learning of qwen 2 models consists of two stages: offline training and online training Yang et al. (2024d). In the online training stage, the model further refines its performance by interacting with real-time feedback. During this phase, Qwen2 samples multiple candidate responses from the current policy, and a reward model evaluates and ranks them, selecting the most and least preferred responses to form preference pairs. These preference pairs are then fed back into the DPO process, allowing the model to iteratively improve its decision-making and instruction-following capabilities with each episode. This real-time refinement further aligns the model with human preferences, ensuring that it delivers more accurate and contextually appropriate responses. By integrating DPO at both stages, Qwen2 not only enhances its alignment with human preferences but also optimizes resource efficiency, making it suitable for handling large-scale models.
@@
The Llama 3 Herd of Models. In Llama3 Herd of Models AI@Meta (2024), post-training procedure includes relatively simple methods such as SFT, reject sampling(RS) and DPO rather than complex reinforcement learning algorithms demanding extra complexity. After the supervised fine-tuning stage, several batches of preference data have been collected using the best performing models from the previous rounds of training. The distribution of the policy model is optimized in each training round based on the collected quality data. The Llama3 team also confirms that compared to on-policy alogorithms like PPO Schulman et al. (2017b), DPO does reduce computational costs and perform better, especially on instruction following benchmarks like IFEval Zhou et al. (2023). To leverage the advantages of DPO, Llama3 team applies two modifications to the original DPO algorithm. Firstly, special formatting tokens including header and terminator are masked out in from both chosen and rejected responses in the loss to stabilize the DPO training, which is based on the observation that these special tokens contributing to loss may induce undesired model behaviors such as tail repetition or abruptly generating termination tokens. Secondly, an additional negative log-likelihood (NLL) loss term is applied to DPO with a scaling coefficient of 0.2 on the chosen sequences, helping further stabilize DPO by maintaining desired formatting for generation and preventing the decrease of log probability of chosen responses Pang et al. (2024).Enhanced Safety DPO approach is a crucial part of ensuring that the model not only follows instructions accurately but also adheres to strict safety guidelines, reducing the likelihood of generating harmful or inappropriate contents.
@@
DeepSeek LLM. DeekSeek LLM DeepSeek-AI et al. (2024) base models are enhanced through post-training using DPO.  DeepSeek-AI et al. (2024) observed an increase in repetitive output as the proportion of math SFT data increased, which they attributed to the repetitive reasoning patterns present in such data. Weaker models often struggle to generalize these patterns effectively, leading to undesirable repetitive responses. To address this issue and enhance the models’ robustness, DPO was employed to further align the SFT models by utilizing a preference-based training approach. This approach leveraged pair-wise comparisons based on criteria such as helpfulness and harmlessness. Upon evaluation, the DPO-trained models demonstrated significant improvements across almost all key metrics, including performance in multilingual tasks such as creative writing, question answering, and instruction following. Moreover, the models also exhibited enhanced mathematical reasoning and logical deduction skills, underscoring the positive impact of the DPO training process in refining language model alignment.
@@
Mixtral of Experts. Mixtral 8x7B Jiang et al. (2024a) is a Sparse Mixture of Experts (SMoE) model with publicly available weights. To improve its ability to follow complex instructions,  Jiang et al. (2024a) applied a two-step training approach. First, they employed SFT on an instruction-focused dataset to enhance model’s responsiveness to diverse commands. This was followed by DPO on a dataset with paired human feedback, allowing for further refinement based on user preferences. After completing the post-training process, Mixtral’s performance significantly exceeded that of well-known models such as GPT-3.5 Turbo OpenAI (2022), Claude-2.1 Anthropic (2024), Gemini Pro Team et al. (2023), and Llama 2 70B Touvron et al. (2023c) in human evaluation benchmarks, highlighting the effectiveness of this combined training methodology.
@@
Baichuan2. Baichuan 2 is a series of large-scale multilingual language models with 7 billion and 13 billion parameters, trained from scratch on a vast dataset of 2.6 trillion tokens Yang et al. (2023). In the post-training stage, Baichuan2 is aligned with limited corpus of human-annotated data to improve the model’s performance with respect to vulnerability safety issues. To be specific,  Yang et al. (2023) is algined with a red-teaming process that incorporates six types of safety attacks and over 100 granular safety value categories. Around 1K annotated data was used for initialization. By leveraging DPO, Baichuan 2 effectively addresses specific vulnerability issues, such as generating harmful or biased content, even with a limited amount of annotated safety-related data. By integrating DPO, Baichuan 2 demonstrates significant improvements in handling complex safety challenges, mitigating risks like the generation of misinformation, biased content, or other undesirable behaviors that can arise in large-scale multilingual models.
@@
Yi: Open Foundation Models. Yi model family AI et al. (2024) encompasses a series of language and multi-modal models. The training of Yi model family proves that delicate fine-tuning framework can effectively save computational costs while maintaining same training results. In typical fine-tuning framework, reference or reward model is used to predict a batch of data or give a scalar reward value after certain action in specific environment. The time and computational resource spent on reference/reward model can not be neglected. In the next step, the target model will use data from previous step to calculate loss and carry out gradient descent in order to update parameters. In the training framework of Yi model family, multi-model scheduling framework is adopted to facilitate support multiple backends for different LLMs in a single job AI et al. (2024). Compared to PPO, DPO algorithm can facilitate the intermediate results from the reference mode. Reusing data from previous stage effectively improves the training speed and save the resources costs.